OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (y – 1) = 3(x + 2) B. (x + 2) = 3(y – 1) C. (x – 2) = 3(y + 1) D. (y + 1) = 3(x – 2) @TheSmartOne @MW3030

TheSmartOne (thesmartone):

ok, so what is point-slope form?

OpenStudy (anonymous):

y - y1 = m(x - x1)

TheSmartOne (thesmartone):

ok, and our slope is 3 plug that in place of m

OpenStudy (anonymous):

y - y1 = 3(x - x1)

TheSmartOne (thesmartone):

now our point is (-2,1) so in place of y_1, we have to put the y value of the point. What is the y-value of the point (-2,1)

OpenStudy (anonymous):

1?

TheSmartOne (thesmartone):

correct, so now we have y - 1 = 3(x-x1)

OpenStudy (anonymous):

Okay

TheSmartOne (thesmartone):

now what is x_1 going to be?

OpenStudy (anonymous):

And then 2 would go in place of the x1 right?

TheSmartOne (thesmartone):

-2** yup :)

OpenStudy (anonymous):

Awesome so it would be D?

TheSmartOne (thesmartone):

nope look at it carefully

TheSmartOne (thesmartone):

y - 1 = 3(x - (-2)) what is x - (-2) = ?

OpenStudy (anonymous):

x + 2

OpenStudy (anonymous):

So A?

TheSmartOne (thesmartone):

correct! :D

OpenStudy (anonymous):

Alrighty! Only 2 more yeah!

TheSmartOne (thesmartone):

Great :D